OpenStudy (anonymous):

The sum of the two roots of a quadratic equation is 5 and their product is -6. What is the quadratic equation?

OpenStudy (jamesj):

Suppose a and b are the roots of a quadratic equation, p(x). Then (one) form of p(x) is p(x) = (x-a)(x-b) = x^2 - (a+b)x + ab Notice that a+b is the sum of the roots and ab is the product. Can you now solve your problem?

OpenStudy (anonymous):

Yes, even I thought of it quite differently.

OpenStudy (anonymous):

\[x + y = 5\]

OpenStudy (anonymous):

\[xy=-6\]

OpenStudy (anonymous):

multiplying he first equation by y

OpenStudy (anonymous):

\[xy + y^2 = 5y\]

OpenStudy (anonymous):

so -6 + y^2 = 5y

OpenStudy (anonymous):

Is that correct?

OpenStudy (jamesj):

Yes, so rearranging: y^2 - 5y - 6 = 0. You can decide which method is shorter.

OpenStudy (anonymous):

I think yours is better hehe.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I did not know of the second method ... thanks!